Gita invests $2642 in a savings account for 3 years that earns 3.2% compounded quarterly.


How much interest will she have earned at the end of 3 years?


1
Expert's answer
2021-04-29T17:46:22-0400

"A=P(1 +\frac {r} {n}) ^{nt}"


"A=\$2,642(1+\frac {0.032}{4})^{4×3}"


"A=\$2,642 (1+0.008)^{12}"


"A=\$2,907.09"


Interest ="\$2,907.09-\$2,642=\$265.09"